# SVDD 2024: The Inaugural Singing Voice Deepfake Detection Challenge

### v1

@@ 3.2 Protocols
Note that the training and development sets available on Zenodo are incomplete because of licensing issues of some bonafide datasets. To fully retrieve the dataset, we asked the participants to first download all the remaining bonafide datasets on their own by agreeing to the terms and conditions and then follow the instructions we provided [^4] https://github.com/SVDDChallenge/CtrSVDD˙Utils. Participants can refer to the statistics in Table 2 as a guide to verify the completion of their downloads and generation.
@@ 3.4 Baseline solutions
We have developed two baseline systems for the challenge: one that uses raw waveforms and another that employs linear frequency cepstral coefficients (LFCCs) as front-end features: The raw waveform system is an AASIST [20]-based system. The LFCC system used 60 coefficients, with a 512 sample window and 160 sample hop size. The LFCC features pass through several downsampling residual convolution blocks and a linear layer connecting it to the graph attention network backend of [20].
@@ 4.1 WildSVDD database
We have continued to gather data annotations from social media platforms following a method similar to the SingFake project [3]. The WildSVDD dataset has been expanded to approximately double the original size of SingFake, now featuring 97 singers with 2007 deepfake and 1216 bonafide songs. The annotators, who are familiar with the singers they cover, manually verified the user-specified labels during the annotation process to ensure accuracy, especially in cases where the singer(s) did not actually perform certain songs. We cross-check the annotations against song titles and descriptions, and manually review any discrepancies for further verification. We have verified the accessibility of all URLs in the dataset as of March 28th and removed any that were inaccessible. The WildSVDD dataset now includes Korean singers, making Korean the third most represented language in the dataset. To help track changes between the SingFake and WildSVDD datasets, we have added a ”SingFake_Set” column that indicates the original partition of an annotation in the SingFake dataset. Annotations that lack a value in this column are new additions to the WildSVDD dataset. We form two test sets: Test A denotes a newly formed testing dataset including new samples, while Test B denotes the hardest test set T04 detailed in [3].
@@ 5.1 CtrSVDD results from the team submissions
Table 5: Overview of the top-8 ranked submission results. The best performance in each column is bolded.
@@ 5.2 Per-attack and per-dataset analysis
For the per-dataset EER, we calculate the EER based on the metadata of the source bonafide datasets, specifically identifying which dataset the target speaker belongs to. For the per-attack EER, A09-A13 is calculated using all bonafide data except ACESinger, along with the deepfake data generated by the corresponding methods. The A14 EER is calculated using the ACESinger bonafide portion and the deepfakes generated by the ACE Studio tool, making it identical to the per-dataset EER for ACESinger. Note that we used a slightly different metric during the challenge but changed post-challenge. Previously, per-attack EERs on A09-A14 were calculated using all bonafide song clips against deepfake clips for each category, which could be what was reported by participants.
The per-attack results in Table 5 do not show a uniform decrease across all attacks as ranking increases, indicating that the best-performing systems may not be the most robust. This observation aligns with findings in speech antispoofing research [23]. A notable example is the performance of A11 and A12. Compared to our baseline systems [8], A12 shows significant improvement while slightly underperforming in other methods’ attacks. In contrast, A14 maintains a considerable performance gap compared to other systems. Furthermore, although the top four systems achieved similar EERs and shared the second-place ranking, their robustness differs significantly. NBU_MISL’s system is considerably less robust to ACESinger, dropping to 19th place when A14 is considered, while the others maintain their top positions.
A similar phenomenon is evident in the per-dataset results. While KiSing and M4Singer datasets yielded overall low EERs, ACESinger consistently produced EERs around 50%, approximating random guessing. The performance gap observed with ACESinger likely stems from inconsistencies between ACESinger and the rest of the CtrSVDD dataset. Notably, two teams achieved 93% and 84% on ACESinger, indicating some ability to establish a decision boundary on this dataset, albeit with incorrect decisions.
Both the per-method and per-dataset EERs highlight the ongoing challenge of generalization to unseen methods and acoustic channels. This underscores the need for increased research efforts focused on developing systems that are more robust to these out-of-distribution conditions.
@@ 5.3 Solution strategies
All the top four teams share similar strategies. Their general schemes are illustrated in Figure 1. The top-ranked team, “Fosafer Speech,” used additional datasets besides CtrSVDD: HiFi-TTS [24], OpenSinger [25], CSD [26], itako-Singing, JSUT-Song, Namine_ritsu_utagoe_db, no7-singing, PJS [27], PoPCS [17], URSing [28], which are all among the list of allowable external training datasets in our evaluation plan [7]. They also generated additional training data, nhv_share, and made it public [^8] https://drive.google.com/file/d/1h36C6mWvywIYXSPErDf2tcYfPuRkGKoQ/  [^9] https://drive.google.com/file/d/1LhbH2-yNe˙ZHcmTxqS5q7XQYlL1˙JM˙-/, one week before the challenge submission deadline as requested. They fused three subsystems and incorporated adversarial training [29] to improve domain adaptation in the third system. The team “NBU_MISL” developed two subsystems focused on low and high frequencies, respectively. Both used feature aggregation from Chinese HuBERT and the corresponding frequency band. They augmented the dataset with HiFi-GAN [30] vocoded audio. The team “I2R-ASTAR” developed five subsystems with various layer aggregation strategies [31] and variants of the RawBoost data augmentation method for each subsystem. The team “Qishan” developed two subsystems with different SSL features. Each subsystem follows a Sensitive Layer Select (SLS) classifier that uses an adaptive weight allocation method [32] to aggregate SSL features and pool the feature map to a score. The score with a larger absolute value is selected for submission.
@@ 6 Discussions and conclusions
The CtrSVDD track of the SVDD challenge was a notable success, attracting 47 submissions, with 36 surpassing the baseline performance. The top teams employed diverse and advanced techniques, such as self-supervised learning, ensemble learning, and adversarial training, demonstrating significant innovation in the field. Detailed system descriptions from eight teams provided valuable insights for future research. This success highlights the progress in deepfake detection for singing voices and sets the stage for further advancements and improvements.

### v2

@@ 3.2 Protocols
Note that the training and development sets available on Zenodo are incomplete because of licensing issues with some bonafide datasets. To fully retrieve the dataset, we asked the participants to first download all the remaining bonafide datasets on their own by agreeing to the terms and conditions and then follow the instructions we provided [^4] https://github.com/SVDDChallenge/CtrSVDD˙Utils. Participants can refer to the statistics in Table 2 as a guide to verify the completion of their downloads and generation.
@@ 3.4 Baseline solutions
We have developed two baseline systems for the challenge: one that uses raw waveforms and another that employs linear frequency cepstral coefficients (LFCCs) as front-end features: The raw waveform system is an AASIST [20]-based system. The LFCC system uses 60 coefficients, with a 512 sample window and 160 sample hop size. The LFCC features pass through several downsampling residual convolution blocks and a linear layer connecting it to the graph attention network backend of [20].
@@ 4.1 WildSVDD database
We gathered data annotations from social media platforms following a method similar to the SingFake project [3]. The WildSVDD dataset has been expanded to approximately double the original size of SingFake, now featuring 97 singers with 2007 deepfake and 1216 bonafide songs. The annotators, who were familiar with the singers they covered, manually verified the user-specified labels during the annotation process to ensure accuracy, especially in cases where the singer(s) did not actually perform certain songs. We cross-checked the annotations against song titles and descriptions, and manually reviewed any discrepancies for further verification. We verified the accessibility of all URLs in the dataset as of March 28th and removed any that were inaccessible. The WildSVDD dataset now includes Korean singers, making Korean the third most represented language in the dataset. To help track changes between the SingFake and WildSVDD datasets, we have added a ”SingFake_Set” column that indicates the original partition of an annotation in the SingFake dataset. Annotations that lack a value in this column are new additions to the WildSVDD dataset. We form two test sets: Test A denotes a newly formed testing dataset including new samples, while Test B denotes the hardest test set T04 detailed in [3].
@@ 5.1 CtrSVDD results from the team submissions
Table 5: Overview of the top-8 ranked submission results. The best performance in each column is bolded and the second is underlined.
@@ 5.2 Per-attack and per-dataset analysis
For the per-dataset EER, we calculate the EER based on the metadata of the source bonafide datasets, specifically identifying which dataset the target speaker belongs to. For the per-attack EER, A09-A13 is calculated using all bonafide data except ACESinger, along with the deepfake data generated by the corresponding methods. The A14 EER is calculated using the ACESinger bonafide portion and the deepfakes generated by the ACE Studio tool, making it identical to the per-dataset EER for ACESinger. Note that we used a slightly different metric during the challenge but changed post-challenge. Previously, per-attack EERs on A09-A14 were calculated using all bonafide song clips against deepfake clips for each category [8], which could be what was reported by participants.
Table 5 provides a detailed analysis of the top eight ranked submissions. The per-attack results do not show a consistent decline across all attacks as rankings improve. Notably, the performance on A11 and A12 suggests that the top-performing systems may not necessarily be more robust to different types of attacks. This observation is consistent with findings in speech anti-spoofing research [23]. Besides, The systems’ performance on A12 is generally worse compared to other attacks, which is consistent with our baseline analysis [8]. Furthermore, although the top four systems achieved similar EERs, their robustness differs significantly. The “NBU_MISL” system is less robust to KiSing and ACESinger compared to other top systems, dropping to 19th place when ACESinger is factored in.
A similar phenomenon is evident in the per-dataset results. The best performance on KiSing was achieved by the team “Star”, which was ranked 8th. While KiSing and M4Singer datasets yield overall low EERs, ACESinger consistently produced EERs around 50%, approximating random guessing. The performance gap observed with ACESinger is likely due to inconsistencies between ACESinger and the rest of the CtrSVDD dataset. Notably, two teams achieved 93% and 84% on ACESinger, indicating some ability to establish a decision boundary on this dataset, albeit with incorrect decisions.
Both the per-attack and per-dataset EERs highlight the ongoing challenge of generalization to unseen generation methods and acoustic channels. This underscores the need for increasing research efforts focused on developing systems that are more robust to these out-of-distribution conditions.
@@ 5.3 Solution strategies
All the top four teams share similar strategies. Their general schemes are illustrated in Figure 1. The top-ranked team, “Fosafer Speech,” used additional datasets besides CtrSVDD: HiFi-TTS [24], OpenSinger [25], CSD [26], itako-Singing, JSUT-Song, Namine_ritsu_utagoe_db, no7-singing, PJS [27], PoPCS [17], URSing [28], which are all among the list of allowable external training datasets in our evaluation plan [7]. They also generated additional training data, nhv_share, and made it public [^8] https://drive.google.com/file/d/1h36C6mWvywIYXSPErDf2tcYfPuRkGKoQ/  [^9] https://drive.google.com/file/d/1LhbH2-yNe˙ZHcmTxqS5q7XQYlL1˙JM˙-/, one week before the challenge submission deadline as requested. They fused three subsystems and incorporated adversarial training [29] to improve domain adaptation in the third system. The team “NBU_MISL” developed two subsystems focused on low and high frequencies, respectively. Both used feature aggregation from Chinese HuBERT and the corresponding frequency band. They augmented the dataset with HiFi-GAN [30] vocoded audio. The team “I2R-ASTAR” developed five subsystems with various layer aggregation strategies [31] and variants of the RawBoost data augmentation method for each subsystem. Please refer to [32] for more details of their system. The team “Qishan” developed two subsystems with different SSL features. Each subsystem follows a Sensitive Layer Select (SLS) classifier that uses an adaptive weight allocation method [33] to aggregate SSL features and pool the feature map to a score. The score with a larger absolute value is selected for submission.
@@ 6 Discussions and conclusions
The CtrSVDD track of the SVDD challenge was a notable success, attracting 47 submissions, with 37 surpassing the baseline performance. The top teams employed diverse and advanced techniques, such as self-supervised learning, ensemble learning, and adversarial training, demonstrating significant innovation in the field. Detailed system descriptions from eight teams provided valuable insights for future research. This success highlights the progress in deepfake detection for singing voices and sets the stage for further advancements and improvements.